# with application to sea and land ice

### v1

@@ Introduction
Non-Newtonian fluid models are often used as continuum representations of complex physical systems which, over suitable spatial and temporal scales, display a fluid-like behavior. Examples of these complex systems include polymer solutions [larson2013], colloidal suspensions [Guazzelli18], granular flows [jop2006], and Earth’s mantle [ranalli1995]. Two additional examples explored in this article are ice sheets (i.e. land ice) [greve2009] and sea ice [feltham2008].
Non-Newtonian models are generally written as a system of partial differential equations (PDEs) that consist of conservation laws and constitutive models. The constitutive model for a fluid, which is also referred to as its rheology, relates inner stresses with the deformation of the material. Traditionally, rheological models have been built in two ways: by finding a function with a simple analytical expression that fits the data (phenomenological models) or by deriving the model from first principles using e.g. kinetic theory [jenkins1983, toppaladoddi2025]. Finding accurate models via these two approaches is often a difficult task due to the inherent complexity of certain materials.
Data-driven methods are an alternative approach to finding accurate constitutive models for complex materials. These methods integrate data with mathematical models to infer physically-consistent material laws. One such class of methods assumes a predefined functional form for the rheology that contains unknown scalar parameters. Typically, gradient-based optimization is used to choose these parameters in order to fit the observation data. In doing so, the governing equations are either enforced as constraints [hu2024, akerson2025, kontogiannis2025], or their residuals are penalized as done in physics-informed neural networks (PiNNs), [wang2025, lardy2025].
Using a predefined functional form for the rheology requires knowledge about the material, which may not be available. Recently, several methods have been proposed to avoid using predefined functional forms through the application of scientific machine learning (sciML) techniques. In [saadat2022, mahmoud2024, thakur2024], for instance, rheologies are identified within a library of functions using sparsification techniques that promote simple models. Alternatively, and of particular importance to this work, a class of methods represent the rheology with neural networks (NNs) [reyes2021, bruer2022, lennon2023, parolini2025, sunol2025, dediego2025]. The availability of powerful machine learning libraries with automatic differentiation (AD) capabilities and the approximation properties of NNs make this approach particularly appealing. Before discussing details of this related work, we summarize our approach.
@@
Next, we review other contributions that also represent rheological models with NNs. They differ in e.g. what kind of data the NNs are trained on (stress or velocity data) or what physical and mathematical principles are enforced in the model. In [lennon2023], viscoelastic rheologies are represented with NNs and trained with experimental polymer data. The NNs are trained on stress data by solving the time-dependent initial value problem governing a viscoelastic rheology. In [parolini2025], convex NNs are fitted to stress data via standard regression to represent the rheology for a set of shear-thinning fluids. Convexity of the NN is enforced to ensure well-posedness of the underlying PDEs. A class of monotone rheologies built with an NN are fitted to both rheology and velocity data in [sunol2025]. This approach is applied to synthetic data to recover known rheological models. Finally, in [dediego2025], the authors of this article apply the method described here to infer an unknown rheology that reproduces synthetic sea ice Lagrangian floe data; we summarize these results briefly in section 4.2 below. This article represents an extension of [dediego2025] in that we derive key physical and mathematical principles that the rheology should satisfy in two-dimensions, we present a streamlined implementation applicable to a wider class of fluids, and we explore the method’s accuracy and sensitivity to data errors in synthetic settings in land and sea ice models for which the rheology is known.
@@
The main contributions of this work are as follows: (1) We derive and discuss the incorporation of three critical physical and mathematical principles for the rheological model: isotropy, the existence of an underlying potential of dissipation, and the second law of thermodynamics. The rheological models satisfying these principles are more general than the class of rheologies considered in [parolini2025, sunol2025]. (2) We treat the governing equation during training using methods from PDE-constrained optimization, which are combined with NN libraries. This should be contrasted with e.g. PiNNs-based methods, as in [thakur2024, wang2025], where the sum of the data misfit and the PDE residual is minimized, which may affect the numerical PDE solution accuracy in the presence of large data errors. (3) We illustrate numerically that our framework performs robustly for velocity and stress data that include large errors. In particular, we infer two known rheological models (Glen’s law for land ice, and the effective shear viscosity of the viscous-plastic model for sea ice) from noisy data. To demonstrate our ability to discover unknown models, we infer a rheology from sea ice data generated with a discrete element method (DEM) that tracks individual ice floes.
The limitations of this work are: (1) We confine ourselves to the inference of a parameter-dependent effective shear viscosity function. This function fully defines the constitutive relation for incompressible models, as used for land ice, while it only partially characterizes this relationship for compressible flows, as utilized in the context of sea ice. (2) The inferred rheology functions involve only a single external parameter (concentration for sea ice, and temperature for land ice). However, our approach easily extends to multiple external parameters. (3) Our implementation relies on an external coupling between a finite element library for solving PDEs and a machine learning library for parameterizing NN functions. This coupling is not straightforward, particularly when it comes to composing the differentiation capabilities of both libraries. One solution to this is to develop PDE solvers in an AD software, as in [sunol2025, alhashim2025], and to rely on AD capabilities also for the governing PDE rather than on adjoints.
@@
for a function $\mathcal{C}:\mathbb{S}_{2}\to\mathbb{S}_{2}$, where $\mathbb{S}^{2}$ denotes the set of symmetric matrices in $\mathbb{R}^{2\times 2}$. The assumptions regarding explicitness and locality are a constraint that simplifies both our theoretical and computational framework. Relaxing this constraint could be an intriguing subject for future research. In the context of our application to sea ice, locality may be the most restricting of the two assumptions, since granular flows are known to exhibit nonlocal behavior [kamrin2012]. For further information on implicit rheologies, we refer the reader to [rajagopal2006].
@@
holds for all $\boldsymbol{\mathrm{O}}\in\mathbb{O}_{2}$. It is possible to show (see [gurtin1982, Section 37]) that $\mathcal{C}$ is isotropic if and only if there exist two scalar functions $\psi_{i}:\mathbb{R}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$ for $i=1,2$ such that
@@
This characterization of isotropic rheologies allows us to narrow down the rheology inference problem to the discovery of two functions $\psi_{1}$ and $\psi_{2}$. The key physical principle of frame-indifference then automatically holds when we represent $\mathcal{C}$ as in (7) [truesdell2000]. We point out that a compressible Newtonian fluid, for which $\boldsymbol{\sigma}$ depends on $\boldsymbol{\mathrm{D}}\boldsymbol{u}$ linearly, is generally written in terms of two scalar parameters, a bulk viscosity $\zeta$ and a shear viscosity $\eta$, as
@@
Since the necessary optimality condition $\langle\mathcal{J}^{\prime}(\boldsymbol{u}),\boldsymbol{v}\rangle=0$ corresponds to the weak form of (1) when $\boldsymbol{\mathrm{D}}\boldsymbol{u}/\boldsymbol{\mathrm{D}}t=0$, the equivalence is established. Moreover, if $\Phi$ is convex, then $\mathcal{J}$ is strictly convex (since we assume $j$ is strictly convex). Thus, if $\mathcal{J}$ is coercive on $V$, we can establish the existence and uniqueness of solutions to (1) whenever $\boldsymbol{\mathrm{D}}\boldsymbol{u}/\boldsymbol{\mathrm{D}}t=0$, see e.g. [evans2010].
@@
for all velocity fields $\boldsymbol{u}$. Inequality (16) is a consequence of the second law of thermodynamics [morro2023], which states that, in an isolated system, entropy must always increase. This translates into the following condition for $\psi_{1}$ and $\psi_{2}$:
@@
Sections 2.3 and 2.4 outline the computational framework for inferring the effective shear viscosity $\psi_{\boldsymbol{\theta}}$. Its implementation combines two Python modules, the machine learning library PyTorch [paszke2019] and the finite element library Firedrake [ham2023]. We use PyTorch to construct the NN-based function $\psi_{\boldsymbol{\theta}}$ and minimize the loss functions $\hat{\mathcal{J}}_{s}(\boldsymbol{\theta})$ and $\hat{\mathcal{J}}_{v}(\boldsymbol{\theta})$. The minimization is carried out with PyTorch’s implementation of the LBFGS algorithm. LBFGS computes an approximation of the Hessian of the loss function with the secant method [wright1999]. The gradients of $\mathcal{J}_{s}$ and of the penalty terms in (34) are calculated using PyTorch’s automatic differentiation module.
Gradients of $\mathcal{J}_{v}$ are computed with the adjoint method. In particular, the adjoint method enables efficient and accurate computations of the gradients of each of the terms inside the sum in $\mathcal{J}_{v}$. For computing the gradients, the boundary value problem (22) and its adjoint equation must be solved. Thus, evaluating the gradient of $\mathcal{J}_{v}$ requires the solution of $2N$ PDEs, with NN of these being nonlinear in general. We solve these NN-based PDEs with the finite element method using Firedrake. In particular, for all one-dimensional PDEs in this paper, we use continuous piecewise-linear elements for the horizontal velocity $u$ on a mesh with $50$ cells of equal length. The required coupling between Firedrake and PyTorch is facilitated by novel functionalities that have recently been implemented in Firedrake [bouziani2024]. These novel enhancements of Firedrake allow us to combine the differentiation capabilities of both modules. We also remark that, in our finite element-based approach, the PDE (22) is a constraint to the velocity loss $\mathcal{J}_{v}$ whose discretization is solved accurately at each optimization step.
@@
Ice sheets and mountain glaciers make up the large masses of ice known as land ice. These bodies of ice play an important role in Earth’s climate in many ways: they regulate sea levels [edwards2021], influence ocean currents [lohmann2021], and carve out topographical features in the landscape [koppes2015], among others. Over long timescales, land ice is usually modeled as a slow-moving incompressible viscous fluid, such that the inertial terms in (25) can be neglected. Under the condition of incompressibility, the rheology is completely defined by the effective shear viscosity function $\psi$, as explained in section 2.2.
Glen’s law is the most common choice of rheological law for land ice [glen1958]. It establishes a power law formula for the effective shear viscosity that depends on the temperature $T$. Using the terminology from section 2.1, we can write Glen’s law as
@@
Here, we explore the recovery of Glen’s law from shear stress and velocity measurements. To generate training data, we use a series of rheometer-like one-dimensional simulations with different parameters to expose different rheology regimes. An alternative to using several one-dimensional simulations to generate data is to use a small number of two-dimensional simulations to collect training data, as done in [sunol2025]. Our setup consists of an infinitely-long slab of ice of uniform thickness $L$ sliding down a bedrock inclined at an angle $\alpha$, see figure 1. The external forces $\boldsymbol{f}$ correspond with gravitational forces. Since we consider all quantities but temperatures to be normalized, we set the thickness of the slab to $L=1$ and write the gravitational acceleration vector as $\boldsymbol{f}=(\sin{\alpha},-\cos{\alpha}).$ This setup corresponds with the one-dimensional configuration examined in section 2.2. Therefore, (1) can be reduced to a single equation for the horizontal velocity $u$:
@@
As discussed in section 2.5, we minimize the loss functions $\hat{\mathcal{J}}_{s}$ and $\hat{\mathcal{J}}_{v}$ with the LBFGS algorithm implemented in PyTorch. Here, we present detailed results on the minimization of the loss functions when inferring Glen’s law from stress and velocity data. The numerical behavior we observe for the problems in section 4 is similar. We focus on the cases with $\sigma_{s}=0.1$ and $\sigma_{v}/u_{\max}=0.1$ for the stress and velocity data noise, respectively. In figure 3, we show the evolution of the loss functions and the $\ell^{\infty}$-norm of the gradients $\nabla_{\boldsymbol{\theta}}\hat{\mathcal{J}}_{s}$ and $\nabla_{\boldsymbol{\theta}}\hat{\mathcal{J}}_{v}$. To motivate the use of the LBFGS method, we also show results obtained with the ADAM algorithm, [adam2014]. To compare the two algorithms fairly, we show the loss functions and their gradient norms against the number of function evaluations, instead of the number of iterations. For ADAM, the number of function evaluations coincides with the number of iterations. However, this is not the case for LBFGS due to the linesearch method used at each iteration. For these computations, the penalty parameters are $\beta_{1}=10^{-5}$ when minimizing $\hat{\mathcal{J}}_{s}$ and $\beta_{1}=5\times 10^{-5}$ when minimizing $\hat{\mathcal{J}}_{v}$. For both loss functions, we choose $\beta_{2}=10^{14}$. For ADAM, we set the learning rate to $0.01$ and, at each iteration, we compute the gradient over the complete set of data, not over a randomized subset. We choose this learning rate value after comparing the method’s performance with four other learning rates. To warm-start the minimization of the velocity loss, we perform 10 LBFGS iterations of the stress loss with noisy data. This results in an initial guess that visibly differs from the exact rheology yet is improved compared to using random NN weights.
@@
A two-dimensional problem such as that presented in figure 4 contains a rich wealth of data and mimics realistic glaciological configurations better than the slab problem in figure 1. For an incompressible fluid, such a two-dimensional simulation could have been used to generate a training dataset. This approach is followed in [sunol2025], where a shear-thinning Carreau-Yasuda model is inferred from a single two-dimensional computation in which the fluid is driven by a pressure gradient through a constriction. Future work could compare the effects of data errors from one and two-dimensional settings on the accuracy of the inference.
@@
Sea ice, which is a critical component of Earth’s climate system, consists of an aggregate of ice segments, called ice floes, driven by ocean and wind drag. It is typically modeled as a continuum when studying its dynamics over spatial scales larger than $100\text{\,}\mathrm{k}\mathrm{m}$ [feltham2008, blockley2020]. Continuum models treat sea ice like a viscous fluid that evolves over a two-dimensional horizontal plane [coon1974]. Momentum balance is expressed with a thickness-averaged version of (1) that can be written as follows:
@@
The viscous-plastic model, also known as Hibler’s model, is the most commonly used model for simulating sea ice dynamics [hibler1979]. Practically all Earth system models use this model or its variants [blockley2020]. The viscous-plastic model treats sea ice like a compressible fluid with a plastic behavior determined by an elliptical yield curve. This rheology can be fully characterized with an equation for the equilibrium pressure $p$ and the two rheological functions $\psi_{1}$ and $\psi_{2}$ introduced in section 2.1. The two rheological functions are given by
@@
Discrete element methods (DEMs) resolve the dynamics of individual ice floes, including floe-scale processes such as collisions, ridging, and fracturing. Considered a promising avenue for sea ice modeling, the high computational costs involved in DEM simulations restricts their use to small scale simulations. For this reason, sea ice modules in Earth System Models rely on continuum models whose accuracy is limited by the use of phenomenological parameterizations to represent complex floe scale phenomena [feltham2008].
To demonstrate the potential of our approach in revealing unknown rheological models from data with substantial errors, we infer an effective shear viscosity that reproduces data generated with SubZero, a sophisticated DEM for sea ice modeling [manucharyan2022]. This DEM models the dynamics of irregularly-shaped polygonal ice floes that interact through collisions, see figure 5. More complex processes, such as fracturing and ridging, are modeled in this DEM. However, we deactivate these features when generating the data used in the inference. The numerical results presented here are similar to those in [dediego2025], and we refer to that work for further information on the rheology inferred from DEM data and several tests proving the rheology’s ability to generalize beyond the training dataset. The results presented here differ from [dediego2025] in that two rheological models are inferred from stress and velocity data independently (in [dediego2025], the model inferred from velocity data uses the model inferred from stress data as an initial guess in the optimization). Moreover, unlike [dediego2025], the NN $\chi$ in (29) is precomposed with a logarithm. We remark that these changes do not result in substantially different models.
@@
Two rheological models inferred from DEM data are presented in figure 7, one computed by minimizing the stress loss $\hat{\mathcal{J}}_{s}$ and the other the velocity loss $\hat{\mathcal{J}}_{v}$. The left panel displays the two stress-strain relationships, together with the DEM data points. The right panel presents the effective shear viscosity $\psi$ of the two models. Velocity profiles computed by solving the training dataset problems with the two models are shown in figure 8. We also include the values of the (unpenalized) loss functions $\mathcal{J}_{s}$ and $\mathcal{J}_{v}$ in table 3. The rheological models predicted with both stress and velocity data reveal a transition from a shear-thickening to a shear-thinning viscosity as the sea ice concentration $A$ increases. The model inferred from stress data achieves a high degree of accuracy in reproducing the velocity profiles of the training dataset (see the red dotted line in figure 8). However, some discrepancies between model and velocity data arise for low ocean velocities $U_{o}$ and $A=0.85$; this is unsurprising, given the substantial errors present in the DEM stress data due to interpolation from Lagrangian to Eulerian data, and the necessary spatial and temporal averaging. The model inferred from velocity data, however, corrects this discrepancy, yielding an impeccable match between model prediction and velocity training data. Further tests presented in [dediego2025] demonstrate the models’ accuracy in reproducing DEM data outside of the training dataset.
@@
In this work, we only consider a single external parameter (temperature for land ice and concentration for sea ice). However, the rheologies of complex systems often depend on additional parameters; for example, the size and thickness of ice floes are known to influence the overall sea ice rheology [wilchinsky2006, feltham2008, herman2022]. Including multiple external parameters increases the size of the training dataset and thus the computational cost of fitting the NN parameters. Since in this work (and most related work cited in the introduction), we assume that our training data results from (possibly expensive) simulations rather than physical observations or measurements, we are able to, at least in principle, generate training data at will. In the presence of several external parameters, one could generate (and use during training) only a small number of selected combinations of external parameters, instead of a tensor grid of parameters as in this work. Which parameter combinations to generate data for could be chosen through random sampling, or systematically such that the simulations expose the rheological behavior in different regimes. Selective data generation that optimally informs the rheology would be an interesting topic of future research. It could be based on sensitivities of the loss function, or other criteria to identify optimal training data motivated from recent advances in active learning or experimental design.

### v2

@@ Introduction
Non-Newtonian fluid models are often used as continuum representations of complex physical systems which, over suitable spatial and temporal scales, display a fluid-like behavior. Examples of these complex systems include polymer solutions [26], colloidal suspensions [14], granular flows [21], and Earth’s mantle [38]. Two additional examples explored in this article are ice sheets (i.e. land ice) [13] and sea ice [11].
Non-Newtonian models are generally written as a system of partial differential equations (PDEs) that consist of conservation laws and constitutive models. The constitutive model for a fluid, which is also referred to as its rheology, relates inner stresses with the deformation of the material. Traditionally, rheological models have been built in two ways: by finding a function with a simple analytical expression that fits the data (phenomenological models) or by deriving the model from first principles using e.g. kinetic theory [20, 43]. Finding accurate models via these two approaches is often a difficult task due to the inherent complexity of certain materials.
Data-driven methods are an alternative approach to finding accurate constitutive models for complex materials. These methods integrate data with mathematical models to infer physically-consistent material laws. One such class of methods assumes a predefined functional form for the rheology that contains unknown scalar parameters. Typically, gradient-based optimization is used to choose these parameters in order to fit the observation data. In doing so, the governing equations are either enforced as constraints [19, 1, 23], or their residuals are penalized as done in physics-informed neural networks (PiNNs), [45, 25].
Using a predefined functional form for the rheology requires knowledge about the material, which may not be available. Recently, several methods have been proposed to avoid using predefined functional forms through the application of scientific machine learning (sciML) techniques. In [40, 30, 42], for instance, rheologies are identified within a library of functions using sparsification techniques that promote simple models. Alternatively, and of particular importance to this work, a class of methods represent the rheology with neural networks (NNs) [39, 6, 28, 34, 41, 8]. The availability of powerful machine learning libraries with automatic differentiation (AD) capabilities and the approximation properties of NNs make this approach particularly appealing. Before discussing details of this related work, we summarize our approach.
@@
Next, we review other contributions that also represent rheological models with NNs. They differ in e.g. what kind of data the NNs are trained on (stress or velocity data) or what physical and mathematical principles are enforced in the model. In [28], viscoelastic rheologies are represented with NNs and trained with experimental polymer data. The NNs are trained on stress data by solving the time-dependent initial value problem governing a viscoelastic rheology. In [34], convex NNs are fitted to stress data via standard regression to represent the rheology for a set of shear-thinning fluids. Convexity of the NN is enforced to ensure well-posedness of the underlying PDEs. A class of monotone rheologies built with an NN are fitted to both rheology and velocity data in [41]. This approach is applied to synthetic data to recover known rheological models. Finally, in [8], the authors of this article apply the method described here to infer an unknown rheology that reproduces synthetic sea ice Lagrangian floe data; we summarize these results briefly in section 4.2 below. This article represents an extension of [8] in that we derive key physical and mathematical principles that the rheology should satisfy in two-dimensions, we present a streamlined implementation applicable to a wider class of fluids, and we explore the method’s accuracy and sensitivity to data errors in synthetic settings in land and sea ice models for which the rheology is known.
@@
The main contributions of this work are as follows: (1) We derive and discuss the incorporation of three critical physical and mathematical principles for the rheological model: isotropy, the existence of an underlying potential of dissipation, and the second law of thermodynamics. The rheological models satisfying these principles are more general than the class of rheologies considered in [34, 41]. (2) We treat the governing equation during training using methods from PDE-constrained optimization, which are combined with NN libraries. This should be contrasted with e.g. PiNNs-based methods, as in [42, 45], where the sum of the data misfit and the PDE residual is minimized, which may affect the numerical PDE solution accuracy in the presence of large data errors. (3) We illustrate numerically that our framework performs robustly for velocity and stress data that include large errors. In particular, we infer two known rheological models (Glen’s law for land ice, and the effective shear viscosity of the viscous-plastic model for sea ice) from noisy data. To demonstrate our ability to discover unknown models, we infer a rheology from sea ice data generated with a discrete element method (DEM) that tracks individual ice floes.
The limitations of this work are: (1) We confine ourselves to the inference of a parameter-dependent effective shear viscosity function. This function fully defines the constitutive relation for incompressible models, as used for land ice, while it only partially characterizes this relationship for compressible flows, as utilized in the context of sea ice. (2) The inferred rheology functions involve only a single external parameter (concentration for sea ice, and temperature for land ice). However, our approach easily extends to multiple external parameters. (3) Our implementation relies on an external coupling between a finite element library for solving PDEs and a machine learning library for parameterizing NN functions. This coupling is not straightforward, particularly when it comes to composing the differentiation capabilities of both libraries. One solution to this is to develop PDE solvers in an AD software, as in [41, 2], and to rely on AD capabilities also for the governing PDE rather than on adjoints.
@@
for a function $\mathcal{C}:\mathbb{S}_{2}\to\mathbb{S}_{2}$, where $\mathbb{S}^{2}$ denotes the set of symmetric matrices in $\mathbb{R}^{2\times 2}$. The assumptions regarding explicitness and locality are a constraint that simplifies both our theoretical and computational framework. Relaxing this constraint could be an intriguing subject for future research. In the context of our application to sea ice, locality may be the most restricting of the two assumptions, since granular flows are known to exhibit nonlocal behavior [22]. For further information on implicit rheologies, we refer the reader to [37].
@@
holds for all $\boldsymbol{\mathrm{O}}\in\mathbb{O}_{2}$. It is possible to show (see [15, Section 37]) that $\mathcal{C}$ is isotropic if and only if there exist two scalar functions $\psi_{i}:\mathbb{R}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$ for $i=1,2$ such that
@@
This characterization of isotropic rheologies allows us to narrow down the rheology inference problem to the discovery of two functions $\psi_{1}$ and $\psi_{2}$. The key physical principle of frame-indifference then automatically holds when we represent $\mathcal{C}$ as in (7) [44]. We point out that a compressible Newtonian fluid, for which $\boldsymbol{\sigma}$ depends on $\boldsymbol{\mathrm{D}}\boldsymbol{u}$ linearly, is generally written in terms of two scalar parameters, a bulk viscosity $\zeta$ and a shear viscosity $\eta$, as
@@
Since the necessary optimality condition $\langle\mathcal{J}^{\prime}(\boldsymbol{u}),\boldsymbol{v}\rangle=0$ corresponds to the weak form of (1) when $\boldsymbol{\mathrm{D}}\boldsymbol{u}/\boldsymbol{\mathrm{D}}t=0$, the equivalence is established. Moreover, if $\Phi$ is convex, then $\mathcal{J}$ is strictly convex (since we assume $j$ is strictly convex). Thus, if $\mathcal{J}$ is coercive on $V$, we can establish the existence and uniqueness of solutions to (1) whenever $\boldsymbol{\mathrm{D}}\boldsymbol{u}/\boldsymbol{\mathrm{D}}t=0$, see e.g. [10].
@@
for all velocity fields $\boldsymbol{u}$. Inequality (16) is a consequence of the second law of thermodynamics [32], which states that, in an isolated system, entropy must always increase. This translates into the following condition for $\psi_{1}$ and $\psi_{2}$:
@@
Sections 2.3 and 2.4 outline the computational framework for inferring the effective shear viscosity $\psi_{\boldsymbol{\theta}}$. Its implementation combines two Python modules, the machine learning library PyTorch [35] and the finite element library Firedrake [16]. We use PyTorch to construct the NN-based function $\psi_{\boldsymbol{\theta}}$ and minimize the loss functions $\hat{\mathcal{J}}_{s}(\boldsymbol{\theta})$ and $\hat{\mathcal{J}}_{v}(\boldsymbol{\theta})$. The minimization is carried out with PyTorch’s implementation of the LBFGS algorithm. LBFGS computes an approximation of the Hessian of the loss function with the secant method [47]. The gradients of $\mathcal{J}_{s}$ and of the penalty terms in (34) are calculated using PyTorch’s automatic differentiation module.
Gradients of $\mathcal{J}_{v}$ are computed with the adjoint method. In particular, the adjoint method enables efficient and accurate computations of the gradients of each of the terms inside the sum in $\mathcal{J}_{v}$. For computing the gradients, the boundary value problem (22) and its adjoint equation must be solved. Thus, evaluating the gradient of $\mathcal{J}_{v}$ requires the solution of $2N$ PDEs, with NN of these being nonlinear in general. We solve these NN-based PDEs with the finite element method using Firedrake. In particular, for all one-dimensional PDEs in this paper, we use continuous piecewise-linear elements for the horizontal velocity $u$ on a mesh with $50$ cells of equal length. The required coupling between Firedrake and PyTorch is facilitated by novel functionalities that have recently been implemented in Firedrake [5]. These novel enhancements of Firedrake allow us to combine the differentiation capabilities of both modules. We also remark that, in our finite element-based approach, the PDE (22) is a constraint to the velocity loss $\mathcal{J}_{v}$ whose discretization is solved accurately at each optimization step.
@@
Ice sheets and mountain glaciers make up the large masses of ice known as land ice. These bodies of ice play an important role in Earth’s climate in many ways: they regulate sea levels [9], influence ocean currents [29], and carve out topographical features in the landscape [24], among others. Over long timescales, land ice is usually modeled as a slow-moving incompressible viscous fluid, such that the inertial terms in (25) can be neglected. Under the condition of incompressibility, the rheology is completely defined by the effective shear viscosity function $\psi$, as explained in section 2.2.
Glen’s law is the most common choice of rheological law for land ice [12]. It establishes a power law formula for the effective shear viscosity that depends on the temperature $T$. Using the terminology from section 2.1, we can write Glen’s law as
@@
Here, we explore the recovery of Glen’s law from shear stress and velocity measurements. To generate training data, we use a series of rheometer-like one-dimensional simulations with different parameters to expose different rheology regimes. An alternative to using several one-dimensional simulations to generate data is to use a small number of two-dimensional simulations to collect training data, as done in [41]. Our setup consists of an infinitely-long slab of ice of uniform thickness $L$ sliding down a bedrock inclined at an angle $\alpha$, see figure 1. The external forces $\boldsymbol{f}$ correspond with gravitational forces. Since we consider all quantities but temperatures to be normalized, we set the thickness of the slab to $L=1$ and write the gravitational acceleration vector as $\boldsymbol{f}=(\sin{\alpha},-\cos{\alpha}).$ This setup corresponds with the one-dimensional configuration examined in section 2.2. Therefore, (1) can be reduced to a single equation for the horizontal velocity $u$:
@@
As discussed in section 2.5, we minimize the loss functions $\hat{\mathcal{J}}_{s}$ and $\hat{\mathcal{J}}_{v}$ with the LBFGS algorithm implemented in PyTorch. Here, we present detailed results on the minimization of the loss functions when inferring Glen’s law from stress and velocity data. The numerical behavior we observe for the problems in section 4 is similar. We focus on the cases with $\sigma_{s}=0.1$ and $\sigma_{v}/u_{\max}=0.1$ for the stress and velocity data noise, respectively. In figure 3, we show the evolution of the loss functions and the $\ell^{\infty}$-norm of the gradients $\nabla_{\boldsymbol{\theta}}\hat{\mathcal{J}}_{s}$ and $\nabla_{\boldsymbol{\theta}}\hat{\mathcal{J}}_{v}$. To motivate the use of the LBFGS method, we also show results obtained with the ADAM algorithm, [33]. To compare the two algorithms fairly, we show the loss functions and their gradient norms against the number of function evaluations, instead of the number of iterations. For ADAM, the number of function evaluations coincides with the number of iterations. However, this is not the case for LBFGS due to the linesearch method used at each iteration. For these computations, the penalty parameters are $\beta_{1}=10^{-5}$ when minimizing $\hat{\mathcal{J}}_{s}$ and $\beta_{1}=5\times 10^{-5}$ when minimizing $\hat{\mathcal{J}}_{v}$. For both loss functions, we choose $\beta_{2}=10^{14}$. For ADAM, we set the learning rate to $0.01$ and, at each iteration, we compute the gradient over the complete set of data, not over a randomized subset. We choose this learning rate value after comparing the method’s performance with four other learning rates. To warm-start the minimization of the velocity loss, we perform 10 LBFGS iterations of the stress loss with noisy data. This results in an initial guess that visibly differs from the exact rheology yet is improved compared to using random NN weights.
@@
A two-dimensional problem such as that presented in figure 4 contains a rich wealth of data and mimics realistic glaciological configurations better than the slab problem in figure 1. For an incompressible fluid, such a two-dimensional simulation could have been used to generate a training dataset. This approach is followed in [41], where a shear-thinning Carreau-Yasuda model is inferred from a single two-dimensional computation in which the fluid is driven by a pressure gradient through a constriction. Future work could compare the effects of data errors from one and two-dimensional settings on the accuracy of the inference.
@@
Sea ice, which is a critical component of Earth’s climate system, consists of an aggregate of ice segments, called ice floes, driven by ocean and wind drag. It is typically modeled as a continuum when studying its dynamics over spatial scales larger than 100 $0\text{\,}\mathrm{k}\mathrm{m} [1$1, 4]. Continuum models treat sea ice like a viscous fluid that evolves over a two-dimensional horizontal plane [7]. Momentum balance is expressed with a thickness-averaged version of (1) that can be written as follows:
@@
The viscous-plastic model, also known as Hibler’s model, is the most commonly used model for simulating sea ice dynamics [18]. Practically all Earth system models use this model or its variants [4]. The viscous-plastic model treats sea ice like a compressible fluid with a plastic behavior determined by an elliptical yield curve. This rheology can be fully characterized with an equation for the equilibrium pressure $p$ and the two rheological functions $\psi_{1}$ and $\psi_{2}$ introduced in section 2.1. The two rheological functions are given by
@@
Discrete element methods (DEMs) resolve the dynamics of individual ice floes, including floe-scale processes such as collisions, ridging, and fracturing. Considered a promising avenue for sea ice modeling, the high computational costs involved in DEM simulations restricts their use to small scale simulations. For this reason, sea ice modules in Earth System Models rely on continuum models whose accuracy is limited by the use of phenomenological parameterizations to represent complex floe scale phenomena [11].
To demonstrate the potential of our approach in revealing unknown rheological models from data with substantial errors, we infer an effective shear viscosity that reproduces data generated with SubZero, a sophisticated DEM for sea ice modeling [31]. This DEM models the dynamics of irregularly-shaped polygonal ice floes that interact through collisions, see figure 5. More complex processes, such as fracturing and ridging, are modeled in this DEM. However, we deactivate these features when generating the data used in the inference. The numerical results presented here are similar to those in [8], and we refer to that work for further information on the rheology inferred from DEM data and several tests proving the rheology’s ability to generalize beyond the training dataset. The results presented here differ from [8] in that two rheological models are inferred from stress and velocity data independently (in [8], the model inferred from velocity data uses the model inferred from stress data as an initial guess in the optimization). Moreover, unlike [8], the NN $\chi$ in (29) is precomposed with a logarithm. We remark that these changes do not result in substantially different models.
@@
Two rheological models inferred from DEM data are presented in figure 7, one computed by minimizing the stress loss $\hat{\mathcal{J}}_{s}$ and the other the velocity loss $\hat{\mathcal{J}}_{v}$. The left panel displays the two stress-strain relationships, together with the DEM data points. The right panel presents the effective shear viscosity $\psi$ of the two models. Velocity profiles computed by solving the training dataset problems with the two models are shown in figure 8. We also include the values of the (unpenalized) loss functions $\mathcal{J}_{s}$ and $\mathcal{J}_{v}$ in table 3. The rheological models predicted with both stress and velocity data reveal a transition from a shear-thickening to a shear-thinning viscosity as the sea ice concentration $A$ increases. The model inferred from stress data achieves a high degree of accuracy in reproducing the velocity profiles of the training dataset (see the red dotted line in figure 8). However, some discrepancies between model and velocity data arise for low ocean velocities $U_{o}$ and $A=0.85$; this is unsurprising, given the substantial errors present in the DEM stress data due to interpolation from Lagrangian to Eulerian data, and the necessary spatial and temporal averaging. The model inferred from velocity data, however, corrects this discrepancy, yielding an impeccable match between model prediction and velocity training data. Further tests presented in [8] demonstrate the models’ accuracy in reproducing DEM data outside of the training dataset.
@@
In this work, we only consider a single external parameter (temperature for land ice and concentration for sea ice). However, the rheologies of complex systems often depend on additional parameters; for example, the size and thickness of ice floes are known to influence the overall sea ice rheology [46, 11, 17]. Including multiple external parameters increases the size of the training dataset and thus the computational cost of fitting the NN parameters. Since in this work (and most related work cited in the introduction), we assume that our training data results from (possibly expensive) simulations rather than physical observations or measurements, we are able to, at least in principle, generate training data at will. In the presence of several external parameters, one could generate (and use during training) only a small number of selected combinations of external parameters, instead of a tensor grid of parameters as in this work. Which parameter combinations to generate data for could be chosen through random sampling, or systematically such that the simulations expose the rheological behavior in different regimes. Selective data generation that optimally informs the rheology would be an interesting topic of future research. It could be based on sensitivities of the loss function, or other criteria to identify optimal training data motivated from recent advances in active learning or experimental design.